Let Y = C+I+G C = 100+1/2(Y-T) T = t0+(t1 Y) I = 75 (a) Suppose...

Let

Y = C+I+G

C = 100+1/2(Y-T)

T = t0+(t1 Y)

I = 75

(a) Suppose that the initial values of the fiscal parameters are G = 100; t0 = 50 and t1 = 1/3 : Find the equilibrium levels of output, disposable income, consumption and the government budget deficit.
Answer #1

AD = ct It G = 100+ 0.5(4-7)+ 75+ 100 = 100+ 0.5(y- to-t, y) + 175 = 275+0.5 [y- 50- 3 Y ] = P75 + 3y - 25 = 850+ & Y Put AD